OpenStudy (anonymous):

3(x+2) ------ 5 =2x-3 X= ?

OpenStudy (anonymous):

x=3???

OpenStudy (studygurl14):

\[\frac{3(x+2)}{5}=2x - 3\] multiple by 5 \[3(x+2) = 10x - 15\] expand \[3x + 6 = 10x - 15\] combine liek terms \[21 = 7x \] didive by 7 \[x = 3\] You are correct!